# RooseBERT: A New Deal For Political Language Modelling

### v1

@@ 6 Conclusion
We introduced RooseBERT, a domain-specific pre-trained language model designed for English political debates. RooseBERT is developed through two approaches: continued pre-training of BERT (RooseBERT-cont) and training from scratch with randomly initialised weights (RooseBERT-scr), in both cased and uncased versions, and across both base and large architectures. To build RooseBERT, we curated a comprehensive dataset including transcripts from US presidential debates, UK Parliament, Australian Parliament, Scottish Parliament, United Nations General Assembly and Security Council, and the European Parliament. Through an extensive evaluation on four downstream tasks (NER, sentiment analysis, argument component detection, and argument relation prediction), we showed that RooseBERT consistently outperforms both general-purpose baselines, other domain-adapted models and LLMs, demonstrating its effectiveness in analysing political discourse. To the best of our knowledge, RooseBERT is the only language model tailored specifically for political debate content. Future work will focus on expanding RooseBERT to other languages, enabling cross-lingual analysis of political discourse and facilitating comparative studies across different political argumentation styles.
7 Acknowledgments

### v2

@@ 6 Conclusion
We introduced RooseBERT, a domain-specific pre-trained language model for political debates in English. RooseBERT is developed through two approaches: continued pre-training of BERT, and training from scratch with randomly initialised weights, in both cased and uncased versions, and across both base and large architectures. To build RooseBERT, we curated a comprehensive dataset including transcripts from US presidential debates, UK, Australian, European and Scottish Parliament, United Nations General Assembly and Security Council. Through an extensive evaluation on four downstream tasks (sentiment analysis, stance detection, argument component detection, and argument relation prediction), we showed that RooseBERT consistently outperforms both general-purpose baselines, other domain-adapted models and LLMs, demonstrating its effectiveness in analysing political discourse. To the best of our knowledge, RooseBERT is the only language model tailored specifically for political debates.
Future work will focus on expanding RooseBERT to other languages, enabling cross-lingual analysis of political discourse and comparative studies across different political argumentation styles.
Limitations
While RooseBERT demonstrates substantial improvements over general-purpose PLMs and related political-domain models, our work is subject to several limitations. Our pre-training corpus, though comprehensive, is restricted to English-language debates and speeches. This excludes a large proportion of political deliberation conducted in other languages, which limits RooseBERT’s applicability in multilingual or cross-lingual settings. Furthermore, despite spanning multiple geopolitical contexts (e.g., US presidential debates, UK and Australian parliaments, UN and EU proceedings), the corpus primarily reflects formal institutional settings. Informal or grassroots political discourse (e.g., social media discussions, activist debates) remains underrepresented.
Training RooseBERT—particularly the from-scratch (SCR) variant—required significant computational resources (multi-GPU setups, tens of hours of training). This limits the accessibility of further experimentation (e.g., larger architectures, multilingual versions) for resource-constrained researchers.
Finally, while RooseBERT can support the study of political discourse, it may also be misused for manipulative purposes such as automated propaganda generation, speaker profiling, or biased framing detection. Responsible deployment requires careful consideration of these risks.
Acknowledgments
@@
Appendix A Experiments with Large Architectures
In addition to the base architecture, RooseBERT was also trained in cont and scr approaches, each in both cased and uncased variants using the large (i.e., BERT-large) configuration, as larger models have been shown to capture more complex linguistic patterns. The large model follows BERT-Large, with 24 layers, 1024 hidden units, and 16 attention heads.
To the best of our knowledge, the current state of the art does not include pre-trained language models for political debates which utilise a large architecture. Therefore, the only possible comparison for RooseBERT-large is its baseline, BERT-large.
Pre-Training.
Table 9 display the perplexity results of the large architectures in both cased and uncased version on an held-out portion of the dataset.
Models
Sentiment Analysis
Accuracy
Stance Detection
Macro F1 Score
Arg. Component Detection
Macro F1 Score
Arg. Relation Prediction
Macro F1 Score
BERT-large-cased 0.528 $\pm$ 0.000 0.316 $\pm$ 0.096 0.469 $\pm$ 0.005 0.576 $\pm$ 0.198
BERT-large-uncased 0.544 $\pm$ 0.036 0.352 $\pm$ 0.032 0.466 $\pm$ 0.002 0.410 $\pm$ 0.283
RooseBERT-large-cont-cased 0.719 $\pm$ 0.007 0.367 $\pm$ 0.117 0.475 $\pm$ 0.002 0.499 $\pm$ 0.264
RooseBERT-large-cont-uncased 0.689 $\pm$ 0.090 0.300 $\pm$ 0.128 0.410 $\pm$ 0.135 0.598 $\pm$ 0.210
RooseBERT-large-scr-cased 0.709 $\pm$ 0.006 0.419 $\pm$ 0.011 0.475 $\pm$ 0.002 0.687 $\pm$ 0.008
RooseBERT-large-scr-uncased 0.715 $\pm$ 0.005 0.412 $\pm$ 0.008 0.479 $\pm$ 0.002 0.695 $\pm$ 0.006
BERT-large
RooseBERT-cont
large
RooseBERT-scr
large
cased 27.505 2.203 2.531
uncased 7.832 2.280 2.558
Table 9: Perplexity of large architectures evaluated on an held out corpus representing the overall distribution of political debates.
The BERT-large model struggles without a specific pre-training on the political terminology with perplexity values that are comparable to the perplexity of the base architecture. On the other side, RooseBERT performs significantly better than BERT. The perplexity for the cont configurstion is again lower than its counterpart scr, in both versions.
Training the large architecture required more computational power with respect to the base architecture, with a total training time of 35 hours for both the cont and scr versions. As a consequence, due to time constraints, it was only possible to train the large versions of RooseBERT for 125K steps. The hyper-parameters used to train the large version of RooseBERT (see Table 10) are the same used to train the base version with the exception of the number of steps.
Hyper-parameters
RooseBERT-cont
large
RooseBERT-scr
large
Learning Rate Decay Linear Linear
Warmup Steps 10,000 10,000
Total Training Steps 125K steps 125K steps
Total Batch Size 2048 2048
Learning Rate $1e^{-4}$ $1e^{-4}$
Weight Decay 0.01 0.01
Adam $\beta_{1}$, $\beta_{2}$, $\epsilon$ 0.9, 0.98, $1e^{-6}$ 0.9, 0.98, $1e^{-6}$
Vocabulary Type BERT Custom
Vocabulary Size
28,997 (cased)
30,523 (uncased)
28,996 (cased)
30,522 (uncased)
Table 10: Hyper-parameters used to pre-train RooseBERT with the large architecture setting.
Fine-Tuning.
Due to constraints arising from the computational power required to train the large architectures, for the fine-tuning of RooseBERT-Large, we fixed the max sequence length to $256$ tokens while leaving other hyper-parameters unaltered. Table 8 reports the results of the fine-tuning of BERT-large and RooseBERT-large on the four downstream tasks and hyper-parameters discussed in Section 4. Each result is averaged over five runs with different seeds.
In this case, we again observe the superiority of the scr configuration over the cont one, with RooseBERT-large-scr-uncased outperforming all other models on three out of four tasks. Notably, in the large architecture, the BERT-Large models appeared more prone to overfitting under certain random seeds, especially in tasks such as sentiment analysis and stance detection, leading to inconsistent or degraded performance.
Performance gains from scaling up to the large architecture were modest when compared to the base model. An exception is relation classification, where RooseBERT-large achieved a Macro F1 score of 0.695, improving over the base model’s score of 0.674. This findings suggests that, in some cases, a well-optimised base model is sufficient to capture most of the relevant patterns in the data, and that increasing model size does not always guarantee substantial improvements, particularly when data complexity or volume does not fully leverage the larger parameter space. Furthermore, these results underscore the difficulty of tasks that rely on datasets of political debates, highlighting the need for PLMs specifically trained on such data, like RooseBERT.